OpenStudy (anonymous):

Someone PLEASE help me. lim x->4 = (1/sqrt(x)-2) - (4/x-4)

jimthompson5910 (jim_thompson5910):

Plug in x = 4 and you'll find that you have \[\large \infty-\infty\] which is an indeterminate form, what does that mean?

OpenStudy (anonymous):

L'hospital!!!!!!

OpenStudy (anonymous):

Don't you have to rewrite the problem?

OpenStudy (anonymous):

What is L'hospital?!

OpenStudy (anonymous):

pretty sure he doesnt apply in this situation, but i think of him every time i see limits

jimthompson5910 (jim_thompson5910):

exactly, you have to write the RHS into the form \[\large \frac{f(x)}{g(x)}\]

jimthompson5910 (jim_thompson5910):

unfortunately it doesn't work in all cases, be nice if it did

OpenStudy (anonymous):

There is an actual answer to this though, not just indeterminate.

jimthompson5910 (jim_thompson5910):

which is why you need to get it into \[\large \frac{f(x)}{g(x)}\] form and apply L'Hospitals rule

OpenStudy (anonymous):

rewrite (1/sqrt(x)-2) - (4/x-4) in the form of (x-4)-(4*(sqrt(x)-2))/(sqrt(x)-2)*(x-4), the apply Hôpitals rule, which is to differentiate both sides and then take the limit